# Surgical Aggregation: Federated Class-Heterogeneous Learning

### v5

@@ Abstract
The release of numerous chest x-ray datasets has spearheaded the development of deep learning models with expert-level performance. However, they have limited interoperability due to class-heterogeneity – a result of inconsistent labeling schemes and partial annotations. Therefore, it is challenging to leverage these datasets in aggregate to train models with a complete representation of abnormalities that may occur within the thorax. In this work, we propose surgical aggregation, a federated learning framework for aggregating knowledge from class-heterogeneous datasets and learn a model that can simultaneously predict the presence of all disease labels present across the datasets. We evaluate our method using simulated and real-world class-heterogeneous datasets across both independent and identically distributed (iid) and non-iid settings. Our results show that surgical aggregation outperforms current methods, has better generalizability, and is a crucial first step towards tackling class-heterogeneity in federated learning to facilitate the development of clinically-useful models using previously non-interoperable chest x-ray datasets.
{IEEEkeywords}
Federated learning, class-heterogeneity, generalizability, client drift, chest x-ray
@@ 1 Introduction
Deep learning has revolutionized the field of radiology in recent years, especially for its utility in screening life-threatening conditions using chest x-rays (CXRs) [1]. As a result, numerous large-scale CXR datasets have been released for the purpose of training high-performing models with expert-level detection of diseases in the thorax [2, 3, 4, 5]. Since each dataset is curated at different sites using different acquisition parameters and image processing techniques, large-scale CXR datasets are heterogeneous and not independently and identically distributed (iid). Moreover, these datasets are class-heterogeneous due to inconsistent labeling schemes and partial annotations. In other words, each dataset focuses only on detecting a subset of findings that could be present in a patient and annotations present in one dataset may not be present in another – limiting their interoperability. For example, the NIH Chest X-ray 14 and CheXpert datasets contain 14 and 13 disease labels respectively, with only seven in common, making it challenging to leverage these datasets in aggregate to train large-scale models [5, 2].
Interoperability between these datasets is crucial as models trained on larger, diverse, multi-institutional datasets yield higher performance and better generalizability [6, 7, 8]. In the context of bias and fairness, models trained on multi-institutional datasets tend to the least biased towards under-served patient populations [9]. Therefore, there is a critically unmet need to address interoperability between large-scale CXR datasets, in order to train clinically-useful models with a complete representation of abnormalities that may occur in the thorax – even with the presence of class-heterogeneity.
@@
Figure 1: An overview of surgical aggregation. Suppose two sites are curating CXR datasets to train models for their own niche applications; the first site works on cardiomegaly and pneumonia, while the second works on effusion and pneumonia. Using surgical aggregation, the central server aggregates ”knowledge” from each site for their respective tasks to learn a global model that can detect all three tasks simultaneously.
In this work, we propose surgical aggregation, a semi-supervised federated learning framework for aggregating knowledge from class-heterogeneous datasets (Fig. 1). In the aforementioned example, surgical aggregation can leverage the NIH and CheXpert datasets in aggregate to train a large-scale model that can simultaneously predict the presence of all 20 unique disease labels across both datasets – while performing as well as models trained individually on both datasets. We validate our framework by conducting a series of empirical experiments with varying levels of class- and data-heterogeneity to evaluate their stability using simulated iid datasets. Finally, we evaluate the efficacy of our framework in real-world settings using two large-scale non-iid datasets. Our main contributions are three-fold:
  1. 1.
    We propose a federated class-heterogeneous framework for training high-performing models that learn a complete representation of disease labels present across class-heterogeneous datasets.
  2. 2.
    Our extensive validation shows that surgical aggregation is stable to class- and data-heterogeneity and outperforms existing approaches.
  3. 3.
    We show that our framework has the potential to facilitate the development of large-scale models using previously non-interoperable datasets.
Our methods are detailed in Section 3. Convergence analysis for surgical aggregation is in Section 3.1.3. The experimental design is detailed in Section 4. Our results are detailed in Section 5 and discussed in Section 6. All our implementations are available on https://github.com/UM2ii/SurgicalAggregation.
2 Related Work
The prevalent approach to leverage knowledge from the myriad of CXR datasets is training models individually on each dataset. While they yield high-performance in isolation, individually trained models often do not generalize well in the real-world due to domain shift [10, 11, 12, 7, 8]. Cohen et al. [6] demonstrated that the poor generalizability of individually trained models stemmed from not just shift in imaging data, but also from shift in labels due to factors like inconsistent labeling schemes and clinical biases. Furthermore, this siloed approach would result in an ever increasing number of models as more datasets get curated in the future for novel tasks (e.g., detection of COVID-19).
Various groups have explored the idea of leveraging multiple datasets in aggregate to yield high-performing and highly-generalizable models [8, 7, 6, 9]. However, aggregating multi-site data in a centralized training setup is impractical due to patient data privacy. Federated learning (FL) was proposed as promising collaborative learning paradigm to mitigate this problem. Consisting of a central server with multiple clients, FL can train a global model on distributed datasets by aggregating ”knowledge” (i.e., model weights) learned by local models at each client using a strategy like FedAvg [13, 14]. It has enabled training of large-scale models using medical imaging data distributed across multiple institutions without sharing sensitive patient data [14, 15, 16]. However, class- and data-heterogeneity pose a significant challenge to FL.
Figure 2: Illustration of client drift in a FL setup in non-iid settings. As client model parameters move towards the client optima, the global model parameters drift away from the global optima for the network.
Due to data-heterogeneity, local models may learn features that differ significantly from each other (Fig. 2). Thus, different strategies have been proposed in literature for aggregating knowledge from non-iid datasets [17, 18, 19, 20, 21]. However, there is a significant gap in literature addressing class-heterogeneity. Most approaches focus on either detecting 1) the presence of an abnormality, or 2) the disease labels shared across all datasets [8, 7, 6, 9]. Furthermore, most work by other groups in leveraging class-heterogeneous datasets to train models has been in a centralized setup [22, 23, 24].
Kulkarni et al. [25] explored a personalized FL (PFL) approach using two class-heterogeneous CXR datasets. However, PFL requires each client to learn a unique model and potentially results in ambiguous predictions from client-to-client and poor generalizability in the real-world [26, 6]. Gong et al. [27] proposed an ensemble attention distillation-based FL approach to train a global model using NIH and CheXpert datasets. However, this approach learns a multi-class classifier, i.e., the model can predict the presence of only one disease – even if other diseases may be present. Furthermore, an additional dataset is required for distillation at the central server. Dong et al. [28] developed a class-heterogeneous FL framework that used local consistency regularization to generate pseudo-labels for non-local classes at a client. However, this approach is dependent on the representation learned by the unsupervised pre-training step to generate low-uncertainty pseudo-labels – a crucial consideration for non-iid settings.
For segmentation, Miao et al. [29] proposed a class-heterogeneous FL framework that corrects local optimization by providing direction of gradient for non-local tasks using a modified cross-entropy loss. While segmentation is a fundamentally different problem than classification, the authors similarly noted a lack of prior literature for federated class-heterogeneous learning frameworks that leverage multiple datasets in aggregate to train a singular multi-task model.
Therefore, our proposed framework is different from prior work in three major aspects. First, our framework is capable of training a singular multi-label global model that can simultaneously predict the presence of more than one disease. Second, our framework is semi-supervised and does not require an additional dataset. Finally and most importantly, our framework does not require clients to know the tasks nor share the same tasks of other clients to participate in the network – a core assumption of most FL techniques [14, 15, 16].
@@ 3.1 Surgical Aggregation
Surgical aggregation is a semi-supervised, model and task-agnostic FL framework for ”stitching”, i.e., selectively aggregating, task-specific weights from all clients to construct a global task block for the global model – even with the presence of class-heterogeneity in the network. Our framework is built upon the premise that our method can leverage partially annotated class-heterogeneous datasets to learn a sufficiently good approximation of a model trained with a hypothetical fully annotated equivalent.
@@ 3.1.1 Problem Setup
Let us consider a FL setup comprising of $K\in\mathbb{N}$ clients (indexed by $k$) training for $T\in\mathbb{N}$ epochs with communication after every $E\in\mathbb{N}$ epochs, resulting in $\lfloor T/E\rfloor$ communication rounds. Furthermore, each client $k$ has training data $\{\mathbf{x}_{k},y_{k}\}$ with $n_{k}$ samples.
Definition 1. We define a multi-label classifier as comprising of: 1) A representation block $F$ consisting of feature extraction layers, and 2) a task block $G$ consisting of fully-connected layers that interpret extracted features to predict classes, such that $\hat{y}=(G\circ F)(\mathbf{w};\mathbf{x})$ for data $\{\mathbf{x},y\}$.
Constraint 1. Let $C$ denote the complete set of classes distributed across the network and $C_{k}$ be the set of classes for client $k$, such that $C_{k}\subseteq C$ and $C=\bigcup_{k=1}^{K}C_{k}$.
Most FL approaches operate under the constraint that all classes are shared by all clients, i.e., $\forall c\in C_{i}$, $c\in C_{j}\ \forall i\neq j$. However, class-heterogeneity necessitates a broader definition:
3.1.2 Algorithm Overview
Figure 3: Illustration of the surgical aggregation framework constructing a singular global task block using task-specific weights from each client in the network with task block aggregation strategy $g$ (e.g., mean).
At the start of each communication round, clients train locally for $E$ epochs before communicating the updated weights to the central server. The central server aggregates the weights of all layers in the representation block using strategy $f$ (e.g., FedAvg). Our framework uses a modified aggregation strategy to construct a singular task block using task-specific weights from each client in the network.
As shown in Fig. 3, if a class is shared across one or more clients, the weights corresponding to the class in the task block are aggregated using strategy $g$. In our implementation, we define $g$ as the mean of task-specific weights. However, if a class is not shared with any other client, the weights remain unchanged. This strategy enables the selective aggregation of weights applied to the input of task block to construct a global task block to predict all $C$ tasks in the network. The surgical aggregation framework is detailed in Algorithm 1.
@@ 3.1.2 Algorithm Overview
Input: $K$ clients indexed by $k$ with local tasks $C_{k}$, total epochs $T$, epochs before communication $E$, model layers indexed by $l$, initialized model parameters $\mathbf{w}_{0,k}^{(l)}$, representation block strategy $f$, task block strategy $g$;
@@
       for each $\text{each client}\ k\ \text{and each layer}\ l$ do
             $\mathbf{w}_{t,k}^{(l)}\leftarrow SGD(\mathbf{w}_{t-1,k}^{(l)})$
@@
[…]
             for each $\text{each client}\ k\ \text{and each layer}\ l$ do
[…]
                         $\mathbf{w}_{t,k}^{(l)}\leftarrow f(\mathbf{w}_{t}^{(l)})$
                   else
                         for each class $\text{each class c}\in C_{k}$ do
                               $(\mathbf{w}_{t,k}^{(l)})_{c}\leftarrow g((\mathbf{w}_{t}^{(l)})_{c})$
@@
                   end if

### v6

# From Isolation to Collaboration: Federated Class-Heterogeneous Learning for Chest X-Ray Classification

@@ Abstract
Federated learning (FL) is a promising paradigm to collaboratively train a global chest x-ray (CXR) classification model using distributed datasets while preserving patient privacy. A significant, yet relatively underexplored, challenge in FL is class-heterogeneity, where clients have different sets of classes. We propose surgical aggregation, a FL method that uses selective aggregation to collaboratively train a global model using distributed, class-heterogeneous datasets. Unlike other methods, our method does not rely on the assumption that clients share the same classes as other clients, know the classes of other clients, or have access to a fully annotated dataset. We evaluate surgical aggregation using class-heterogeneous CXR datasets across IID and non-IID settings. Our results show that our method outperforms current methods and has better generalizability.
keywords:
Federated learning, Multi-label classification, Class-heterogeneity, Chest X-ray
Data and Code Availability
All code and data used in this study is publicly available. The NIH Chest X-ray 14 dataset (https://nihcc.app.box.com/v/ChestXray-NIHCC) is open access. The CheXpert (https://stanfordmlgroup.github.io/competitions/chexpert/) and MIMIC-CXR-JPG (https://physionet.org/content/mimic-cxr-jpg/2.1.0/) datasets are credentialed access. Our code is available at https://github.com/BioIntelligence-Lab/SurgicalAggregation.
Institutional Review Board (IRB)
This retrospective study used publicly available datasets and was acknowledged by our IRB as exempt.
@@ 1 Introduction
Federated learning (FL) has emerged as a promising way to collaboratively train deep learning (DL) models for chest x-ray (CXR) classification using private datasets distributed across different institutions (Chowdhury et al., 2022; Sheller et al., 2020). Consisting of a central server and clients, FL can collaboratively train a global model by aggregating the weights of local models communicated by each client using a strategy (e.g., FedAvg) while preserving patient privacy (McMahan et al., 2017).
A fundamental challenge in FL is heterogeneity, where differences in data distribution among clients in a network impacts model optimization (Li et al., 2020). This is relevant in medical imaging because each institution uses different acquisition parameters and image processing techniques, resulting in CXR datasets that are heterogeneous and not independently and identically distributed (IID). Moreover, each institution specializes in different tasks due to the costly and time-consuming process of manual annotation, resulting in partially annotated, class-heterogeneous datasets (Miao et al., 2023). In other words, annotations present in one dataset may not be present in another. For example, the NIH Chest X-Ray 14 and CheXpert datasets contain 14 and 13 disease labels respectively, with only seven in common (Wang et al., 2017; Irvin et al., 2019).
Despite its prevalence in medical imaging, class-heterogeneity is a relatively underexplored topic in FL literature. While several groups have proposed techniques to tackle heterogeneity, they assume that all clients share the same set of classes (Li et al., 2021; Li and Wang, 2019; Karimireddy et al., 2020). However, in the real world, this may not always be the case. While some groups have attempted to solve class-heterogeneity, their techniques assume that clients know the classes of other clients (Miao et al., 2023; Dong et al., 2022) or use knowledge distillation with a public dataset in the central server that is either fully annotated (Gudur and Perepu, 2021) or unlabeled (Gong et al., 2021, 2022). Others avoid this challenge using personalized FL (PFL) techniques with client-specific classification heads (Kanhere et al., 2024; Kulkarni et al., 2022).
@@
Figure 1: An overview of surgical aggregation. Suppose two sites want to collaboratively train a CXR classification model using FL. The first site works on cardiomegaly and pneumonia, while the second works on effusion and pneumonia. Using surgical aggregation, the central server can learn a global model to predict all three labels.
[…]
@@ 2.1 Surgical Aggregation
Surgical aggregation is based on the selective aggregation of weights communicated by clients to dynamically build a global classification head that predicts all classes available in the network from distributed, class-heterogeneous datasets. Our method does not rely on clients sharing the same set of classes or handling other non-local classes to cover three types of partial annotations: 1) Classes shared by all clients, 2) Classes shared by some clients, and 3) Classes unique to a client.
@@ 2.1.1 Problem Setup
Let us consider a FL setup comprising of $K\in\mathbb{N}$ clients (indexed by $k$) training for $T\in\mathbb{N}$ epochs with communication after every $E\in\mathbb{N}$ epochs, resulting in $\lfloor T/E\rfloor$ communication rounds. Let $C_{k}$ be the set of local classes for client $k$. We define $C$ as the complete set of classes distributed across the network, such that $C=\bigcup_{k=1}^{K}C_{k}$.
Figure 2: Illustration of surgical aggregation dynamically building a global classification head using class-specific weights.
We define a multi-label classifier $(G\circ F)$ as comprising of: 1) feature extractor $F$, and 2) classification head $G$. Here, $F$ can be any model backbone (e.g., DenseNet) and $G$ is a fully-connected layer.
Suppose each client $k$ has local model $(G\circ F)_{k}$, where $F_{k}$ is the local feature extractor and $G_{k}$ is the local classification head. At an epoch $t$, we denote the weights of the local feature extractor as $\mathbf{w}_{t,k}^{(F)}$ and global feature extractor as $\mathbf{w}_{t}^{(F)}$.
Similarly, we denote the weights of the local classification head as $\mathbf{w}_{t,k}^{(G)}\in\mathbb{R}^{N\times M_{k}}$ and global classification head as $\mathbf{w}_{t}^{(G)}\in\mathbb{R}^{N\times M}$. Here, $N$ is the number of input neurons to the fully-connected layer, $M$ and $M_{k}$ are the cardinality of $C$ and $C_{k}$ respectively (such that $M\geq M_{k}\ \forall k$). We define the class-specific weights of class $c\in C_{k}$ for client $k$ as column vector of $c$ in $\mathbf{w}_{t,k}^{(G)}$ and is denoted $(\mathbf{w}_{t,k}^{(G)})_{c}\in\mathbb{R}^{N}$. If $c\not\in C_{k}$, then $(\mathbf{w}_{t,k}^{(G)})_{c}=\mathbf{0}$.
2.1.2 Overview of Surgical Aggregation
At the start of each communication round, clients train locally using gradient descent for $E$ epochs on local classes before communicating the updated weights to the central server. To update the global model, we use a modified server update strategy comprising of feature extractor update and classification head update steps.
Feature Extractor Update: We aggregate the weights of local feature extractors $\mathbf{w}_{t,k}^{(F)}$ using model aggregation strategy $f$. Here, $f$ can be any model aggregation strategy (e.g., FedAvg). For example, if $f$ is FedAvg, then the weights of the global feature extractor are defined as (McMahan et al., 2017):
$\mathbf{w}_{t}^{(F)}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t,k}^{(l)}$ (1)
Figure 3: Detailed illustration of surgical aggregation building the weights of the global classification head.
[…]
For class $c\in C$ present in $K_{c}\in\mathbb{N}$ clients (such that $K_{c}\leq K$), the global class-specific weights are computed by the mean of all local class-specific weights:
$(\mathbf{w}_{t}^{(G)})_{c}=\frac{1}{K_{c}}\sum_{k=1}^{K}(\mathbf{w}_{t,k}^{(G)})_{c}$ (2)
If the class $c$ is unique to a client (i.e. $K_{c}=1$), Equation 2 says that the global class-specific weights are equal to the local class-specific weights:
$(\mathbf{w}_{t}^{(G)})_{c}=(\mathbf{w}_{t,k}^{(G)})_{c}$ (3)
[…]
At the end of the communication round, the aggregated $\mathbf{w}_{t}^{(F)}$ is communicated back to the clients. However, our method does not directly communicate the aggregated $\mathbf{w}_{t}^{(G)}$ back to the clients because it contains class-specific weights for all classes, including non-local classes $c\not\in C_{k}$ for a client $k$. Instead, our method reconstructs and communicates $\mathbf{w}_{t,k}^{(G)}$ to each clients $k$, using the updated $(\mathbf{w}_{t}^{(G)})_{c}$ for all $c\in C_{k}$.
@@ 2.1.2 Overview of Surgical Aggregation
Input: $K$ clients indexed by $k$ with local classes $C_{k}$, total epochs $T$, epochs before communication $E$, initialized model parameters $\mathbf{w}_{0,k}$, model aggregation strategy $f$
@@
       // client update
       for each $\text{each client}\ k$ do
             $\mathbf{w}_{t,k}\leftarrow SGD(\mathbf{w}_{t-1,k})$
@@
      // server update
       if $\text{mod}(t,E)=0$ then
             // feature extractor update
             $\mathbf{w}_{t}^{(F)}\leftarrow f(\{\mathbf{w}_{t,k}^{(F)}\ \text{: }\forall k\})$
            // classification head update
             for each $\text{each class}\ c\in C$ do
                   $(\mathbf{w}_{t}^{(G)})_{c}\leftarrow mean(\{(\mathbf{w}_{t,k}^{(G)})_{c}\ \text{if }c\in C_{k}\ \text{: }\forall k\})$
@@
            // communicate to clients
             for each $\text{each client}\ k$ do
                   $\mathbf{w}_{t,k}^{(F)}\leftarrow\mathbf{w}_{t}^{(F)}$
                   for each $\text{each class}\ c\in C_{k}$ do
                         $(\mathbf{w}_{t,k}^{(G)})_{c}\leftarrow(\mathbf{w}_{t}^{(G)})_{c}$
                   end for